# Revisiting Intermediate-Layer Matching in Knowledge Distillation: Layer-Selection Strategy Doesn’t Matter (Much)

### v1

@@ Abstract
Knowledge distillation (KD) is a popular method of transferring knowledge from a large “teacher” model to a small “student” model. KD can be divided into two categories: prediction matching and intermediate-layer matching. We explore an intriguing phenomenon: layer-selection strategy does not matter (much) in intermediate-layer matching. In this paper, we show that seemingly nonsensical matching strategies such as matching the teacher’s layers in reverse still result in surprisingly good student performance. We provide an interpretation for this phenomenon by examining the angles between teacher layers viewed from the student’s perspective.
@@ 1 Introduction
Large language models have achieved impressive performance in various NLP tasks (Brown et al.,, 2020; Devlin et al.,, 2019). However, they need a large number of parameters, making the models cumbersome and difficult to run in resource-restricted scenarios. Knowledge distillation (KD; Hinton et al.,, 2015) is a widely adopted method to reduce model parameters by training a small “student” model from a large “teacher.” With KD, the student is often able to retain most of the teacher’s performance while using a fraction of the its parameters (Sun et al.,, 2020).
@@
Common KD approaches can be generally divided into two categories: prediction matching and intermediate-layer matching. Matching the prediction is usually mandatory, as it informs the student of the task to solve. This can be achieved by minimizing the divergence of predicted distributions Hinton et al., (2015); Wen et al., (2023) or using reinforcement learning Li et al., (2024).
@@
Intermediate-layer matching distills the teacher’s hidden states (i.e., intermediate layers) to the student Sun et al., (2019); Jiao et al., (2020); Wang et al., (2021). This approach often involves minimizing the distance between the student’s and teacher’s hidden states (usually with a linear mapping if the dimensions do not match). Since the student model is often shallower than the teacher, a layer-selection strategy is required to specify which teacher layer is matched to each student layer.
Recently, researchers have explored various layer-selection strategies. Sun et al., (2019) match the student’s layers to evenly spaced teacher layers; Passban et al., (2021) learn an attention mechanism over the teacher’s layers; Haidar et al., (2022) match the student’s layers to randomly selected layers from the teacher, in order; and Wang et al., (2021) matches the last student layer to a teacher layer close to the end. Overall, there lacks consensus on the best strategy for layer selection, and different strategies often result in unexpectedly similar performance. For example, Sun et al., (2019) reports roughly 0.5 points of difference in accuracy between different layer-selection strategies, and Jiao et al., (2020) reports roughly 1-2 points difference in accuracy [^1] Evaluated on GLUE dev set on MNLI-m/mm and MRPC, excluding CoLA since it is highly sensitive..
In this work, we observe an intriguing phenomenon that the layer-selection strategy does not affect intermediate-layer matching for KD (much). Surprisingly, even matching teacher layers to the student in reverse order yields similar performance to forward matching. However, we do see that intermediate-layer matching (regardless of the layer-selection strategy) helps KD, compared with no intermediate-layer matching.
@@
In addition, we provide an interpretation for this interesting finding: from the student’s point of view, the angles between two teacher layers are often acute; thus, matching an arbitrary teacher layer pulls the student layer to a similar direction. As a result, intermediate-layer matching indeed benefits KD, although the matching strategy does not matter (much).
We conducted experiments with four matching strategies (forward, reverse, random, and all-to-one) on six datasets (four classification, two generation), where we explored various settings, including different depths and parameter initializations. Our results consistently demonstrate the aforementioned phenomenon; we also performed in-depth analysis, verifying our interpretation.
@@ 2 Background and Related Work
Knowledge Distillation (KD) is a method of transferring rich knowledge contained in a teacher model to a student model. To inform the student of the task, it is essential to match the student’s and teacher’s predictions. For the teacher distribution $p$ and student distribution $q_{\bm{\theta}_{s}}$, Hinton et al., (2015) suggest minimizing the Kullback–Leibler (KL) divergence between them:
$\mathcal{L}_{\text{KL}}(\bm{\theta}_{s})=\mathbb{E}_{\mathrm{y}\sim p(\mathrm{y}|\mathrm{x})}\big{[}\log{\tfrac{p(\mathrm{y}|\mathrm{x})}{q_{\bm{\theta}_{s}}(\mathrm{y}|\mathrm{x})}}\big{]}$ (1)
@@
Other than minimizing KL, different prediction matching approaches have been proposed. When the teacher distribution is diverse, for example, the reverse KL divergence Tu et al., (2020); Gu et al., (2024) is used due to its mode-seeking behavior, i.e., the student only focuses on one of the high-probability regions in the teacher distribution Bishop, (2006). Wen et al., (2023) propose an $f$-divergence KD framework, where symemtric divergences (such as Jensen–Shannon and total variation distance) provide a balance between mode averaging and mode seeking. Reinforcement learning can also be applied to KD Hao et al., (2022); Li et al., (2024), which makes the student aware of its prefix and addresses the exposure bias problem Bengio et al., (2015).
Regarding intermediate-layer matching, it distills the teacher’s hidden states, thus providing additional supervisory signals to the student Sun et al., (2019). Let =$\mathcal{M}=\{(\varsigma_{i},\tau_{i})\}_{i}$ be the mapping between student and teacher layers, i.e., the $\varsigma_{i}$th layer of the student is mapped to the $\tau_{i}$th layer of the teacher. Intermediate-layer matching typically penalizes the distance between the matched layers, given by
$\mathcal{L}_{\text{hid}}(\bm{\theta}_{s},\{\bm{A}_{i}\}_{i})=\sum\nolimits_{i}\operatorname{dist}(\bm{A}_{i}\bm{h}_{\varsigma_{i}}^{(s)},\bm{h}_{\tau_{i}}^{(t)})$ (2)
where $\operatorname{dist}$ is a distance metric (such as mean squared error). The trainable linear operator $\bm{A}_{i}$ transformers the student’s hidden state $\bm{h}_{\varsigma_{i}}^{(s)}$ to the space of the teacher’s hidden state $\bm{h}_{\tau_{i}}^{(t)}$, when their dimensions do not match. Otherwise, $\bm{A}_{i}$ may be an identity matrix.
Intermediate-layer matching can be applied to different types of representations. Traditionally, this is achieved by matching the student’s and teacher’s activations Sun et al., (2019); Sanh, (2019). Other studies match attention logits Jiao et al., (2020), attention’s query–key–value relations Wang et al., (2021), and cross-sample relations Park et al., (2019); Huang et al., (2023). In our work, we focus on matching activations because it is the most fundamental approach in intermediate-layer matching.
Various layer-selection strategies have been proposed for matching a shallow student to a deep teacher. Sun et al., (2019) and Jiao et al., (2020) suggest mapping evenly spaced teacher layers to the student. Passban et al., (2021) match each student layer to a weighted combination of all teacher layers to retain more knowledge. Haidar et al., (2022) randomly reselect a sequence of teacher layers to match with the student after each epoch, so that the student is exposed to different teacher layers.
Overall, different layer-selection strategies perform unexpectedly similarly (as mentioned in §1), which inspires our work. We observe an intriguing phenomenon that the layer-selection strategy does not matter (much), even with unusual mappings such as reverse order; we also provide an interpretation for this phenomenon.
@@
Model Layer Matching #
MNLI-m/mm
@@
Randomly initialized Random 7 71.2/71.2 82.4/58.8 64.4 82.9 32.67 16.70
@@
Student Weights copied Random 12 79.3/78.3 87.5/67.2 82.6 90.7 48.18 22.04
Table 1: Main results on various layer-selection strategies.
3 Approaches and Setups
In this section, we begin by outlining the layer-selection strategies. We then describe the experimental setups, including datasets, metrics, and neural network hyperparameters.
3.1 Layer-Selection Strategies
@@
Intermediate-layer matching requires a strategy to select which teacher layers are matched with which student layers. In this study, we examine the following layer-selection strategies.
@@
Forward Matching. In this variant, lower student layers are matched to lower teacher layers. In particular, we follow Sun et al., (2019) and select evenly spaced teacher layers for matching.
@@
All-to-One Matching. In this variant, all student layers are matched to the middle teacher layer. While matching to one layer is inspired by previous studies Wang et al., (2020, 2021), we slightly modify their approaches (i.e. matching all student layers instead of one), for fair comparison with the rest of our settings.
@@
Reverse Matching. We propose a counterintuitive strategy, where matching is in reverse order (i.e., lower student layers matched to upper teacher layers). This seemingly nonsensical strategy sheds light on the mechanism of intermediate-layer matching.
@@
Random Matching. We choose the same teacher layers as forward matching, then randomly shuffle the order. The order is maintained during distillation. We average the performance across five seeds to evaluate the effect of different random mappings.
@@
Note that the intermediate-layer matching loss is combined with the predictor’s KL loss by $\mathcal{L}=\mathcal{L}_{\text{KL}}+\lambda\mathcal{L}_{\text{hid}}$, where $\lambda$ is a hyperparameter to balance the losses. In addition, we compare the above strategies the No Matching baseline, which disables intermediate-layer matching. In other words, only KL loss is involved in the KD process.
3.2 Datasets and Models

### v2

@@ Abstract
Knowledge distillation (KD) is a popular method of transferring knowledge from a large “teacher” model to a small “student” model. Previous work has explored various layer-selection strategies (e.g., forward matching and in-order random matching) for intermediate-layer matching in KD, where a student layer is forced to resemble a certain teacher layer. In this work, we revisit such layer-selection strategies and observe an intriguing phenomenon that layer-selection strategy does not matter (much) in intermediate-layer matching—even seemingly nonsensical matching strategies such as reverse matching still result in surprisingly good student performance. We provide an interpretation for this phenomenon by examining the angles between teacher layers viewed from the student’s perspective. Our work sheds light on KD practice, as layer-selection strategies may not be the main focus of KD system design, and vanilla forward matching works well in most setups. [^1] The code is released at the following repo: https://github.com/MANGA-UOFA/Layer-Selection
Revisiting Intermediate-Layer Matching in Knowledge Distillation:
Layer-Selection Strategy Doesn’t Matter (Much)
Zony Yu,  Yuqiao Wen,  Lili Mou Dept. Computing Science & Alberta Machine Intelligence Institute (Amii), University of Alberta Canada CIFAR AI Chair, Amii zony249@gmail.com, yq.when@gmail.com, doublepower.mou@gmail.com
@@ 1 Introduction
Large language models have achieved impressive performance in various NLP tasks (gpt4o; deepseek). However, they need a large number of parameters, making the models cumbersome and difficult to run in resource-restricted scenarios. Knowledge distillation (KD; hintonkd, hintonkd) is a widely adopted method to reduce model parameters by training a small “student” model from a large “teacher.” With KD, the student is often able to retain most of the teacher’s performance while using a fraction of its parameters (minillm; deepseek; wen-etal-2025-knowledge; qwen3).
@@
Common KD approaches can be generally divided into two categories: prediction matching and intermediate-layer matching. Matching the prediction is usually mandatory, as it informs the student of the task to solve. This can be achieved by minimizing the divergence of predicted distributions hintonkd; fdistill; sinkhorn or using reinforcement learning hao2022teacher; llmr.
@@
Intermediate-layer matching distills the teacher’s hidden states (i.e., intermediate layers) to the student patientkd; tinybert; ineffectiveness. This approach often involves minimizing the distance between the student’s and teacher’s hidden states (usually with a linear mapping if the dimensions do not match). Also, a layer-selection strategy is required to specify which teacher layer is matched to which student layer.
Traditionally, researchers have explored various layer-selection strategies. patientkd match the student’s layers to evenly spaced teacher layers; alpkd and universal learn an attention mechanism over the teacher’s layers; railkd match the student’s layers to randomly selected layers from the teacher, albeit in sorted order; and minilmv2 match the last student layer to a teacher layer close to the end.
Despite numerous previous efforts, we observe that the effectiveness of existing layer-selection strategies is not convincing, as previous studies often lack proper controlled comparisons. For example, alpkd only compare their work with the vanilla layer-matching baseline, and minilmv2 compare layer-selection strategies without controlling weight initializations. In addition, existing work tends to omit important distillation settings such as decoder-only model architectures. As a result, there is a lack of understanding about how different strategies compare against each other.
In this work, we aim to better understand how layer-selection strategies affect student performance in KD. We conduct systematic controlled experiments across four models (BERT, BART, T5, and Qwen3) on eight tasks (including both classification and generation). We also consider two weight initialization strategies: random initialization and weight copying from the teacher. Our experiments reveal intriguing findings that
  1. 1.
    The layer-selection strategy doesn’t matter (much), as different strategies perform similarly to each other, and
  2. 2.
    Regardless of layer-selection strategies, intermediate-layer matching itself is a highly effective KD method, outperforming the no-matching baseline.
@@
We provide an interpretation for this phenomenon based on geometric analysis of the hidden states: from the student’s point of view, the angles between two teacher layers are often acute; thus, matching any teacher layer pulls the student layer in a similar direction. As a result, intermediate-layer matching indeed benefits KD, although the matching strategy does not matter (much).
Our study offers a practical suggestion for KD: we recommend KD practitioners to focus on other aspects of KD systems (e.g., distillation temperature, choice of $f$-divergence loss functions), while simply using forward matching for intermediate-layer matching as a default strategy, if computing resources are limited.
@@ 2 Background and Related Work
Knowledge Distillation (KD) is a method of transferring rich knowledge contained in a teacher model to a student model. To inform the student of the task, it is essential to match the student’s and teacher’s predictions (bucila). hintonkd suggest minimizing the Kullback–Leibler (KL) divergence between the teacher and student distributions, denoted by $p$ and $q_{\bm{\theta}_{s}}$, respectively. The KL distillation loss is given by
$\mathcal{L}_{\text{KL}}(\bm{\theta}_{s})=\mathbb{E}_{\mathrm{y}\sim p(\mathrm{y}|\mathrm{x})}\Big[\log{\dfrac{p(\mathrm{y}|\mathrm{x})}{q_{\bm{\theta}_{s}}(\mathrm{y}|\mathrm{x})}}\Big]$ (1)
@@
Other than minimizing KL, different prediction matching approaches have been proposed. When the teacher distribution is diverse, for example, the reverse KL divergence engine; minillm is used due to its mode-seeking behavior, i.e., the student only focuses on one of the high-probability regions in the teacher distribution mode-seeking. fdistill propose an $f$-divergence KD framework, where symmetric divergences (such as Jensen–Shannon and total variation distance) provide a balance between mode averaging and mode seeking. Reinforcement learning can also be applied to KD hao2022teacher; llmr, which makes the student aware of its prefix and addresses the exposure bias problem exp-bias.
Regarding intermediate-layer matching, it distills the teacher’s hidden states, thus providing additional supervisory signals to the student patientkd; wen-etal-2025-knowledge. Let $\mathcal{M}=\{(\varsigma_{i},\tau_{i})\}_{i}$ be the mapping between certain student and teacher layers, i.e., the $\varsigma_{i}$th layer of the student is mapped to the $\tau_{i}$th layer of the teacher. Intermediate-layer matching typically penalizes the distance between the matched layers, given by
$\mathcal{L}_{\text{hid}}(\bm{\theta}_{s},\{\bm{A}_{i}\}_{i})=\sum_{i}\operatorname{dist}(\bm{A}_{i}\bm{h}_{\varsigma_{i}}^{(s)},\bm{h}_{\tau_{i}}^{(t)})$ (2)
where $\operatorname{dist}$ is a distance metric (such as mean squared error). The trainable linear operator $\bm{A}_{i}$ transforms the student’s hidden state $\bm{h}_{\varsigma_{i}}^{(s)}$ to the space of the teacher’s hidden state $\bm{h}_{\tau_{i}}^{(t)}$, if their dimensions do not match. Otherwise, $\bm{A}_{i}$ may be an identity matrix.
Intermediate-layer matching can be applied to different types of representations. Traditionally, this is achieved by matching the student’s and teacher’s activations patientkd; distilBERT. Other studies match attention logits tinybert, query–key–value relations minilmv2, and cross-sample relations rkd; fcd. In our work, we focus on matching activations because it is the most fundamental approach in intermediate-layer matching.
Various layer-selection strategies have been proposed for matching a shallow student to a deep teacher. patientkd and tinybert suggest mapping evenly spaced teacher layers to the student. alpkd and universal match each student layer to a weighted combination of all teacher layers to retain more knowledge. railkd randomly reselect a sequence of teacher layers to match with the student (in order) after each epoch, so that the student is exposed to different teacher layers. minilmv2 and ilm-overfitting suggest mapping the last non-prediction student layer to a teacher layer close to the end.
@@
Model
Layer
Matching
MNLI-m/$m$
@@
Randomly
initialized
Out-of-order rand 7 70.5 / 70.5 82.4 / 58.8 64.4 82.9 32.73 16.71
@@
Student Weights copied Out-of-order rand 12 79.0 / 78.0 87.5 / 67.2 82.6 90.7 47.99 22.01
Table 1: Main results. We use BERT on classification tasks, BART on DART, and T5 on WMT16.
Model Layer Matching #
HellaSwag
Acc
CoQA
Exact Match
Teacher 1 63.11 75.02
Student None 2 35.16 33.73
Forward 3 37.85 35.40
Reverse 4 35.01 37.67
All-to-one 5 34.99 37.35
Out-of-order rand 6 35.47 34.80
Table 2: KD results on more challenging tasks using the recent Qwen3 model.
Overall, although previously proposed methods work well within their environments, it remains unclear how various layer-selection strategies compare under a controlled setup, or the extent they contribute to KD. Thus, we address this by conducting systematic investigations across numerous tasks, models, and initialization strategies, and provide an interpretation for our observations.
3 Approaches
@@
Intermediate-layer matching requires a strategy to select which teacher layers are matched with which student layers. In this study, we explore both standard and seemingly nonsensical layer-selection strategies to uncover its effect on intermediate-layer matching KD.
@@
Forward Matching. In this variant, lower student layers are matched to lower teacher layers. In particular, we follow patientkd and select evenly spaced teacher layers for matching.
@@
All-to-One Matching. In this variant, all student layers are matched to the middle teacher layer. While the idea of matching to one layer is proposed by previous studies minilm; minilmv2, we slightly modify their approaches (i.e., matching all student layers instead of one), for fair comparison with the rest of our settings.
@@
Reverse Matching. We experiment with a counterintuitive strategy, where matching is in reverse order (i.e., lower student layers matched to upper teacher layers). This seemingly nonsensical strategy sheds light on the mechanism of intermediate-layer matching.
@@
Out-of-Order Random Matching. We choose the same teacher layers as forward matching, then randomly shuffle the order. The order is maintained during distillation. We average the performance across five seeds to evaluate the effect of different random mappings.
@@
Note that the intermediate-layer matching loss is combined with the predictor’s KL loss by $\mathcal{L}=\mathcal{L}_{\text{KL}}+\lambda\mathcal{L}_{\text{hid}}$, where $\lambda$ is a hyperparameter to balance the losses. We experimentally determined $\lambda=3$ to be a good balance between KL and intermediate-layer matching losses, since it produces improved performance over the No Matching ($\lambda=0$) baseline, while higher values of $\lambda$ may negatively impact performance.
We emphasize that across all layer-selection strategies, we fix the set of teacher layers that are available for the student to learn from. In other words, the only difference among the layer-selection strategies is the order of matching.
4 Results and Analysis
Figure 1: (a) Illustration of the angle calculation. Cosine similarities are shown for (b) MNLI classification, (c) Encoder in the WMT task, (d) and Decoder in WMT (bottom). Orange refers to the setup of random parameter initialization and blue refers to student weights initialized by the teacher.
Setups. We analyzed different layer-selection strategies on both classification and generation tasks. For classification, we adopt the widely used MNLI (mnli), QQP, [^2] https://www.kaggle.com/c/quora-question-pairs QNLI (squad), and SST-2 (sst2) using the BERT model (bert). We also include more challenging tasks, namely HellaSwag (hellaswag) and CommonsenseQA (CoQA; commonsense), using the more recent Qwen3 models (qwen3). For generation, we use the popular DART (dart) and WMT16 En–Ro (wmt16) datasets. We evaluate BART (bart) on the DART benchmark and T5 (t5) on WMT16, respectively. Our selection of tasks and models give a comprehensive coverage of different tasks and model architectures.
For moderate-sized student models (BERT, BART, and T5), we explore two parameter initialization strategies: copying the weights from select teacher layers and random initialization. The former is a practical method used to quickly transfer knowledge from select teacher layers to the student (distilBERT; predistill), and the latter is used to study the effects of layer-selection strategy in isolation. For the Qwen3 large language model, we only perform weight-copying because it is often not feasible to directly finetune large models on downstream tasks without warming up the weights (commonly achieved through pretraining).
More experimental details can be found in Appendix A.
Main Results. In Table 1, we present the main results of our layer-selection experiments. In Rows 1–2, our finetuned teachers perform similarly to previous work bert; dart; fdistill, showing that we have successfully set up the environment for KD experiments.
We examine different layer-selection strategies. As shown in Rows 9–12, the student model achieves similar results across different strategies, with only 2–3 points difference in accuracy for classification tasks and 1–2 points difference in BLEU for generation tasks. Notice that Reverse Matching and Out-of-Order Random Matching appear nonsensical, when in fact they still achieve close performance to Forward Matching, largely outperforming No Matching. The results show that layer-selection strategy has an unexpectedly small effect on student performance.
It should be emphasized that intermediate-layer matching indeed helps KD compared with No Matching, [^3] One exception is the DART experiment with randomly initialized weights, for which we suspect intermediate-layer matching causes the student to overfit. That said, different strategies still perform similarly to conventional Forward Matching, and thus, it does not contradict our general finding. even though the matching strategy does not play a significant role. On MNLI, for example, all strategies improve upon No Matching by six to ten points with random weight initialization and two points with weight-copying.
Next, we take a closer look at how different layer-selection strategies behave under the two parameter initialization settings. To reiterate, weight-copying is a simple and practical method of transferring the teacher’s knowledge to the student (distilBERT; predistill). We also experiment with randomly initialized students in order to disentangle the effects of layer-matching. In Rows 4–7, we see that layer-selection strategies perform similarly to one another, and for the most part better than No Matching.
Results on Challenging Tasks. We further compare different layer-selection strategies on the more challenging HellaSwag and CoQA tasks, which involve more reasoning. As such, we employ the newer Qwen3 models. As shown in Table 2, we observe a similar trend as in our main results, i.e., different strategies generally produce similar results. This suggests the generality of our findings.
An Interpretation Based on Vector Angles. A curious question arises from these observations: why does intermediate-layer matching help KD, but different layer-selection strategies perform similarly? To answer this, we measure the angles between the teacher’s layers, viewed from the student. Specifically, we measure the angles formed by two teacher layers’ and one student layer’s vector representations, depicted in Figure 1a. We show the phenomenon in the MNLI and WMT16 En–Ro datasets in Figures. 1b, 1c and 1d. We see that in both randomly initialized and weight-copied settings, the cosine similarity is positive, suggesting that the angles are mostly acute. In other words, the student layer is pulled to the same general direction regardless of which teacher layer it is matched to. This finding is consistent with shortgpt and ineffectiveness, where they show that different layers may contain similar knowledge.
Additional Results. We further analyze the student depth in Appendix B and experimental stability in Appendix C.
5 Concluding Remarks
In this paper, we observe an intriguing phenomenon that, although intermediate-layer matching helps knowledge distillation, the layer-selection strategy does not matter (much); we also provide an interpretation based on the angles of teacher and student layers. Our work suggests potential limitations and oversights in previous work, where researchers present various heuristic layer matching methods when training their distilled systems, but their effect is not comprehensively studied. We advise the KD practitioners to focus their efforts on other areas of KD, for example, loss functions, initializations, and representation learning.
6 Limitations
In our work, we have experimented with various setups, including eight tasks (six classification and two generation), four model architectures, and two parameter initialization methods. Although the results are generally consistent, there is one exception that intermediate-layer matching does not help in the DART setup. This is understandable as empirical findings are often noisy. We suspect that it is due to the student model overfitting to the teacher’s representations, since we are training a wider student model (compared to T5 and BERT students) from randomly initialized weights on a small dataset. That said, this does not contradict our conclusions, as all layer-selection strategies still perform equally bad.
Additionally, we clarify that our work focuses on KD in the fine-tuning regime (for a certain task) instead of pretraining, due to the limited resources. It is noticed that fine-tuning is the typical setting of how most practitioners perform KD. We also mention that this paper is intended for an NLP audience, therefore non-NLP experiments (such as computer vision) are outside the scope of this work.
It is also worth mentioning that our work does not suggest intermediate-layer matching is unhelpful for KD. Rather, we present an interesting phenomenon that the layer-selection strategy plays an insignificant role in the process. We argue that future studies on layer selection should have a more rigorous comparison of its effect.
7 Acknowledgments
We thank the reviewers and area chairs for their efforts. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Amii Fellow Program, the Canada CIFAR AI Chair Program, a donation from DeepMind, and the Digital Research Alliance of Canada (alliancecan.ca).
Model Depth
Layer Matching
MNLI-m/$m$
Acc
DART
BLEU
WMT16
BLEU
Teacher 12-layer 1 84.5 / 84.1 48.80 25.90
Student 3-layer None 2 77.4 / 76.5 46.32 22.36
Forward 3 79.7 / 78.8 47.94 22.65
Reverse 4 79.2 / 78.2 48.45 21.57
All-to-one 5 79.4 / 78.7 47.10 21.89
Out-of-order rand 6 79.0 / 78.0 48.18 22.04
6-layer None 7 82.1 / 81.3 46.88 24.91
Forward 8 83.5 / 82.9 48.45 25.00
Reverse 9 82.1 / 80.9 48.45 24.30
All-to-one 10 82.3 / 81.8 48.39 24.44
Out-of-order rand 11 82.3 / 81.5 48.03 24.38
9-layer None 12 84.2 / 83.3 46.05 25.88
Forward 13 84.1 / 83.4 47.66 25.67
Reverse 14 83.2 / 82.4 47.01 25.11
All-to-one 15 83.2 / 82.5 46.95 25.43
Out-of-order rand 16 84.4 / 83.3 47.37 25.41
Table 3: Performance of different layer-selection strategies on students of different depths. Student’s parameters are initialized by copying the weights of the teacher.
Model Run
MNLI-m/$m$
Acc
DART
BLEU
WMT16
BLEU
3-layer
Student
Randomly
Initialized
1 71.2 / 71.2 32.44 16.05
2 72.2 / 71.8 32.41 16.90
3 70.8 / 71.1 33.33 16.95
4 70.5 / 70.8 33.13 17.01
5 67.9 / 67.8 32.35 16.65
Mean
70.5±1.4
/70.5±1.4
32.73±0.41 16.71±0.35
Weights
Copied
1 79.3 / 78.3 48.18 21.79
2 78.5 / 77.4 48.49 21.93
3 79.7 / 78.6 47.65 21.86
4 79.2 / 78.5 48.08 22.53
5 78.5 / 77.3 47.54 21.95
Mean
79.0±0.47
/78.0±0.56
47.99±0.35 22.01±0.27
Table 4: Out-of-Order Random Matching experiments on MNLI, DART, and WMT16 En–Ro. For each task and parameter initialization strategy, we computed the mean and standard deviation of five runs.
Appendix A Datasets and Models